# Quantum field theory treatment of neutrino spin-flavor oscillations in a magnetic field

### v1

@@ 6 Quantum mechanical description of spin-flavor oscillations
In this section, we analyze the quantum mechanical description of spin-flavor oscillations of Majorana neutrinos in an external magnetic field. This study can be done in both flavor and mass bases. We take the flavor basis since it is more common in the standard quantum mechanical approach.
@@
We suppose that two flavor neutrinos, $\nu_{e}$ and $\nu_{\mu}$, propagate along the $z$-axis and interact with the transverse external magnetic field $\mathbf{B}=(B,0,0)$, as in Sec. 4. Despite one can reveal the nature of neutrinos only in the mass basis, we can still assume that $\nu_{\lambda}$, $\lambda=e,\mu$, are Majorana particles. The interaction of these neutrinos with the magnetic field is owing to the transition magnetic moment $\mu$. The effective Schrodinger equation in the basis $\nu^{\mathrm{T}}=(\nu_{e},\nu_{\mu},\bar{\nu}_{e},\bar{\nu}_{\mu})$ was obtained in Ref. [7],
@@
where $E_{\nu}$ is the mean energy of neutrinos, and $\theta$ is the vacuum mixing angle defined in Eq. 2.1. We already mentioned in Sec. 2 (see also Ref. [21]) that the magnetic moments matrix $(\mu_{ab})=\mu\sigma_{2}$ is hermitian and purely imaginary in the mass basis. Taking into account Eq. 2.1, we get that this matrix in the flavor basis reads $(M_{\lambda\lambda^{\prime}})=U(\mu_{ab})U^{\dagger}=\mu\sigma_{2}$, that is used in Eq. (6.1).
@@
Now, we can compare the quantum mechanical and quantum field theory approaches to neutrino spin-flavor oscillations. If both neutrinos and charged leptons, which are used to produce and detect neutrinos, are ultrareletivistic, we can put $E_{\nu}=E$, $\mathfrak{E}_{\nu}=\mathfrak{E}$, and $t\approx L$ in Eq. (6.23). Then, we omit the small correction $\delta P_{\nu_{e}\to\bar{\nu}_{\mu}}$ in Eq. (5.17). In this situation, $P_{\nu_{e}\to\bar{\nu}_{\mu}}^{(0)}$ in Eq. (5.16) coincides with the transition probability in Eq. (6.23), i.e. both approaches provide the same description of spin-flavor oscillations.
7 Conclusion
In the present work, we have developed the QFT description of neutrino spin-flavor oscillations in a magnetic field. We were based on the QFT formalism, which was originally developed in Refs. [13, 14, 15], where neutrinos are virtual particles. This method was used in Ref. [13] to study neutrino-to-antineutrino oscillations in vacuum. If one deals with Majorana neutrinos with transition magnetic moments, $\nu\leftrightarrow\bar{\nu}$ oscillations are analogous to spin-flavor transitions in a magnetic field. We have adapted the results of Ref. [13] for the description of spin-flavor oscillations.
@@ 7 Conclusion
The basics of the QFT approach for neutrino spin-flavor oscillations have been developed in Sec. 3. We have obtained the matrix element corresponding to the transition $\nu_{\beta}\to\bar{\nu}_{\alpha}$ in an external magnetic field. The dressed propagators for neutrino mass eigenstates, accounting for the magnetic interaction, have been derived in Sec. 4 for neutrinos with arbitrary energies. Then, we have considered the case of ultrarelativistic particles. Note that we have obtained both diagonal and nondiagonal dressed propagators in Sec. 4.
@@
Finally, we have derived the transition probability for $\nu_{e}\to\bar{\nu}_{\mu}$ oscillations in a magnetic field in Sec. 5 within the QFT approach. This transition probability has the leading term, which was shown in Sec. 6 to coincide with the result of the standard quantum mechanical approach. The correction to the leading term, arising from the QFT description, has been also obtained in Sec. 5; cf. Eq. (5.17). Unfortunately, this correction is vanishing for any oscillations channels and any realistic magnetic fields.
The main goal of the present work is methodological. It is aimed to demonstrate that neutrino spin-flavor oscillations can be described purely with help of the standard QFT methods if reasonable approximation are made. The major assumption is the ultrarelativity of charged leptons which are used to produce and detect virtual neutrinos. The next important assumption, which was also noticed in Ref. [19], is the consideration of two mass eigenstates only. This feature allows one to avoid the branching of the Feynman graphs in Fig. 2.
@@
We also make an interesting observation about neutrino oscillations in external fields. The neutrino interaction with background matter or with a magnetic field can be rather trivially incorporated to the neutrino vacuum oscillations description. Indeed, one just adds a linear term with a certain external field to the vacuum effective Hamiltonian for flavor neutrinos. However, from the point of view of QFT, the effect of external fields on neutrino oscillations is an essentially nonperturbative phenomenon. The correct form of the transition probability is obtained only if the infinite number of terms in Eq. (4) is summed up.
@@ Appendix A Propagators of Weyl neutrinos in vacuum
[…]
@@
$S_{a}^{(\xi)}(x-y)=\theta(x_{0}-y_{0})\left\langle 0|\xi(x)\xi^{\dagger}(y)|0\right\rangle-\theta(y_{0}-x_{0})\left\langle 0|\xi^{*}(y)\xi^{\mathrm{T}}(x)|0\right\rangle.$ (A.8)

### v2

@@ 6 Quantum mechanical description of the spin-flavor precession
In this section, we analyze the quantum mechanical description of the spin-flavor precession of Majorana neutrinos in an external magnetic field. This study can be done in both flavor and mass bases. We take the flavor basis since it is more common in the standard quantum mechanical approach.
@@
We suppose that two flavor neutrinos, $\nu_{e}$ and $\nu_{\mu}$, propagate along the $z$-axis and interact with the transverse external magnetic field $\mathbf{B}=(B,0,0)$, as in Sec. 4. Despite one can reveal the nature of neutrinos only in the mass basis, we can still assume that $\nu_{\lambda}$, $\lambda=e,\mu$, are Majorana particles. The interaction of these neutrinos with the magnetic field is owing to the transition magnetic moment $\mu$. The effective Schrödinger equation in the basis $\nu^{\mathrm{T}}=(\nu_{e},\nu_{\mu},\bar{\nu}_{e},\bar{\nu}_{\mu})$ was obtained in Ref. [7],
@@
where $E_{\nu}$ is the mean energy of neutrinos, and $\theta$ is the vacuum mixing angle defined in Eq. (2.1). We already mentioned in Sec. 2 (see also Ref. [22]) that the magnetic moments matrix $(\mu_{ab})=\mu\sigma_{2}$ is hermitian and purely imaginary in the mass basis. Taking into account Eq. (2.1), we get that this matrix in the flavor basis reads $(M_{\lambda\lambda^{\prime}})=U(\mu_{ab})U^{\dagger}=\mu\sigma_{2}$, that is used in Eq. (6.1).
@@
Now, we can compare the quantum mechanical and quantum field theory approaches to the neutrino spin-flavor precession. If neutrinos are ultrareletivistic, we can put $E_{\nu}=E$, $\mathfrak{E}_{\nu}=\mathfrak{E}$, and $t\approx L$ in Eq. (6.23). Then, we omit the small correction $P_{\mathrm{max}}^{(\mathrm{qft})}$ in Eq. (5.16). Under these approximations, the transition probability in Eqs. (5.15) and (5.16) coincides with that in Eq. (6.23), i.e. both approaches provide the same description of the spin-flavor precession.
7 Discussion
In the present work, we have developed the QFT description of neutrino the spin-flavor precession in a magnetic field. We were based on the QFT formalism, which was originally developed in Refs. [14, 15, 16], where neutrinos are virtual particles. This method was used in Ref. [14] to study neutrino-to-antineutrino oscillations in vacuum. If one deals with Majorana neutrinos with transition magnetic moments, $\nu\leftrightarrow\bar{\nu}$ oscillations are analogous to spin-flavor transitions in a magnetic field. We have adapted the results of Ref. [14] for the description of the spin-flavor precession.
@@ 7 Discussion
The basics of the QFT approach for the neutrino spin-flavor precession have been discussed in Sec. 3. We have obtained the matrix element corresponding to the transition $\nu_{\beta}\to\bar{\nu}_{\alpha}$ in an external magnetic field. The dressed propagators for the neutrino mass eigenstates, accounting for the magnetic interaction, have been derived in Sec. 4 for neutrinos with arbitrary energies. Then, we have considered the case of ultrarelativistic particles.
@@
Finally, we have derived the transition probability for the $\nu_{e}\to\bar{\nu}_{\mu}$ precession in a magnetic field in Sec. 5 within the QFT approach. This transition probability has the leading term, which was shown in Sec. 6 to coincide with the result of the standard quantum mechanical approach. Moreover, we have obtained that the dressed propagator in Eq. (4.5) acquires the terms (i) and (ii) which have the QFT nature. These terms cannot be neglected before the transition probability calculation. We have demonstrated in Sec. 5 that the QFT terms in the propagator result in the correction to the leading term in the transition probability which is small if neutrinos are ultrarelativistic; cf. Eqs. (5.15) and (5.16).
In Sec. 5.1, we have discussed a realistic situation when charged leptons have arbitrary energies. The forward scattering approximation is not valid in this case. We have derived the effective transition probability in Eq. (5.21) which depends on the velocities of incoming and outgoing charged leptons. It turns out that $P^{(\mathrm{eff})}_{\nu_{e}\to\bar{\nu}_{\mu}}<P^{(\mathrm{qm})}_{\nu_{e}\to\bar{\nu}_{\mu}}$.
One of the main goals of the present work is methodological. It is aimed to demonstrate that the neutrino spin-flavor precession can be described purely with help of the standard QFT methods if reasonable approximation are made.
The major assumption is the ultrarelativity of particles involved in the system. First, in Sec. 5, we assume that charged leptons are utrarelativistic, i.e. $E_{\alpha,\beta}\gg\mathfrak{m}_{\alpha,\beta}$. The assumption of ultrarelativistic charged leptons was used to derive the transition probability for the spin-flavor precession in the forward scattering approximation. We considered a general situation of charged leptons with arbitrary energies in Sec. 5.1. Thus, the assumption of ultrarelativistic charged leptons is not crucial for the results of the work.
The most important approximation is the ultrarelativity of virtual neutrinos. It can be expressed in the form, $E=(E_{\alpha}+E_{\beta})/2\gg m_{1,2}$. Along with the spin-flavor precession under the influence of a magnetic field, for Majorana particles, one has neutrino-to-antineutrino transitions owing to a nonzero neutrino mass. Thus, the series in Eq. (4), would have a more complicated form. Moreover, the bare propagators, derived in Appendix, have rather simple form for ultrarelativistic neutrinos. Finally, without the assumption of the neutrinos ultrarelativity, the computation of the contour integrals, $I_{+}$ and $I_{-}$, would be impossible in Sec. 5. For neutrinos with arbitrary energies, the equation $E-\mathcal{E}_{\pm}(\mathbf{q})+\mathrm{i}0=0$, which defines the poles, has eight roots which should found analytically. Thus, the assumption of ultrarelativistic neutrinos in crucial to treat the spin-flavor precession in frames of QFT.
The next important approximation, which was also mentioned in Ref. [20], is the consideration of two mass eigenstates only. This feature allows one to avoid the branching of the Feynman graphs in Fig. 2.
In deriving of Eq. (5), we made an assumption of the great propagation distance $L$. It allows one to avoid terms $\propto L^{-2}$ in Eq. (5). It is, however, known that the coherence in neutrino oscillations is lost if $L>L_{\mathrm{coh}}$, where $L_{\mathrm{coh}}$ is the coherence length. It is found that $L_{\mathrm{coh}}\propto L_{\mathrm{osc}}\sim\mathfrak{E}^{-1}$ (see, e.g., Ref. [12]).
In fact, numerous factors can contribute to the coherence length. For instance, it depends on a particular experiment where neutrino oscillations are studied (see, e.g., Ref. [26]). A careful determination of $L_{\mathrm{coh}}$ requires the replacement of the plane waves wavefunctions of charged leptons with wave packets of a finite width $\sigma_{x}$ [15, 26]. This approach is applicable for the studies of neutrino oscillations in vacuum in frames of QFT. In that situation, the expression for the matrix element can be derived analytically for plane waves wavefunctions of charged leptons, with both neutrinos and charged leptons having arbitrary energies.
In our case, when we study neutrinos interacting with an external magnetic field, the integrations in Sec. 5 are quite nontrivial even for plane waves charged leptons. As we mentioned earlier, the matrix element can be derived only in the approximation of ultrarelativistic neutrinos. Therefore, a careful consideration of the decoherence in the spin-flavor precession in frames of QFT is a quite complicated task which is beyond the scope of the present work. We can just roughly estimate the coherence length as $L_{\mathrm{coh}}\sim\sigma_{x}E/\mathfrak{E}$ for ultrarelativistic neutrinos [15]. Another estimate for the coherence length in a magnetic field reads [27] $L_{\mathrm{coh}}\sim\sigma_{x}E^{3}/\mathfrak{E}m_{a}^{2}$.
@@
We also make an interesting observation about neutrino oscillations in external fields. The neutrino interaction with background matter or with a magnetic field can be rather trivially incorporated to the neutrino vacuum oscillations description. Indeed, one just adds a linear term with a certain external field to the vacuum effective Hamiltonian for flavor neutrinos. However, from the point of view of QFT, the effect of external fields on neutrino oscillations is an essentially nonperturbative phenomenon. The correct form of the transition probability is obtained only if all terms in the infinite series in Eq. (4) are summed up.
@@ Appendix A Propagators of Weyl neutrinos in vacuum
where $E_{a}=\sqrt{\mathbf{p}^{2}+m_{a}^{2}}$ and $\mathrm{i}0$ stays for a small imaginary term. We can decompose Eq. (A.1) as $S_{a}^{(\eta)}=S_{a\mathrm{L}}+\tilde{S}_{a\mathrm{R}}$, where
@@
$S_{a}^{(\xi)}(x-y)=\theta(x_{0}-y_{0})\left\langle 0\right|\xi(x)\xi^{\dagger}(y)\left|0\right\rangle-\theta(y_{0}-x_{0})\left\langle 0\right|\xi^{*}(y)\xi^{\mathrm{T}}(x)\left|0\right\rangle.$ (A.8)